OpenStudy (anonymous):

An Algebra 2 student solved the equation 2x + 4 = 5 using the following steps. Is the student's work correct? If not, explain the mistake. 2x + 4 = 5 log25 = x + 4 log 2/log 5 = x + 4 log 2/log 5 - 4 = x x ≈ -3.569

OpenStudy (anonymous):

Wow! man what level is this maths :P

OpenStudy (anonymous):

Lol its algebra 2 xD

OpenStudy (anonymous):

and why is that genius taking logs @Crushshroom

OpenStudy (anonymous):

I don't understand where each step starts and ends, and I don't understand your logs. What are the bases?

OpenStudy (anonymous):

Logs and Exponents confuse me /.| hold on, im going back to the question

OpenStudy (anonymous):

Bases is 10 @MaxwellSmart !

OpenStudy (anonymous):

If no base is mentioned we assume it to be 10

OpenStudy (anonymous):

\(2x + 4 = 5\) What is the first step after the equation?

OpenStudy (anonymous):

-4 right?

OpenStudy (anonymous):

on both sides?

OpenStudy (anonymous):

Is this the first step of the student's solution? \(\log25 = x + 4 \)

OpenStudy (anonymous):

its 2^x+4=5

OpenStudy (anonymous):

@MaxwellSmart do you want to solve this

OpenStudy (anonymous):

I give up. I don't understand the steps you wrote above. You need to use the equation editor or the draw tool.

OpenStudy (anonymous):

@Crushshroom i'll do it for ya!

OpenStudy (anonymous):

This is a waste of time to just figure out what the problem is.

OpenStudy (anonymous):

Thank you ( : I appreciate the help

OpenStudy (anonymous):

Okay well im sorry, first time using this site -.-

OpenStudy (anonymous):

So u take logs on both sides first! log2^x+4= log5 using the power rule of logs x+4= log5/log2 x= 2.321....-4 x= -1.67....

OpenStudy (anonymous):

Then do what I suggested. Use the equation editor or the draw tool. You can also take a pic of your text and attach it. After much discussion, you finally wrote the equation is 2^x + 4 = 5 I have a feeling that is still not it, and it is 2^(x + 4) = 5 or \(\large 2^{x + 4} = 5\)

OpenStudy (anonymous):

YES! I just figured out how to use te equation thing so ill write the next step

OpenStudy (anonymous):

\[\log_{2}5=x+4 \]

OpenStudy (anonymous):

log 2/log 5 = x + 4

OpenStudy (anonymous):

log 2/log 5 - 4 = x

OpenStudy (anonymous):

Hey @MaxwellSmart I think u are the right person to help him!

OpenStudy (anonymous):

x ≈ -3.569

OpenStudy (anonymous):

thats the answer and im a girl xD

OpenStudy (anonymous):

If you guys cant figure it out its okay ^.^

OpenStudy (anonymous):

I think this is what the problem states were the steps the student did. \(\large 2^{x + 4} = 5\) \(\large log_2{5} = x + 4 \) \(\large \dfrac{\log 2}{\log 5} = x + 4\) \(\large \dfrac{\log 2}{\log 5} - 4 = x\) \(\large x \approx -3.569\)

OpenStudy (anonymous):

It is rather hard

OpenStudy (anonymous):

Yes :D

OpenStudy (anonymous):

Omg youre amazing ( :

OpenStudy (anonymous):

Ok, it only took half an hour to figure out what the problem is.

OpenStudy (anonymous):

Let me copy that and number the steps. Then we'll be able to discuss them.

OpenStudy (anonymous):

1. \(\large 2^{x + 4} = 5\) 2. \(\large log_2{5} = x + 4 \) 3. \(\large \dfrac{\log 2}{\log 5} = x + 4\) 4. \(\large \dfrac{\log 2}{\log 5} - 4 = x\) 5. \(\large x \approx -3.569\)

OpenStudy (anonymous):

In step 3 what did u do @MaxwellSmart

OpenStudy (anonymous):

The above is not my solution. The above is a copy if your problem showing the steps the student did. Step 1 is the equation. Step 2 is taking log base two of both sides. It is correct.

OpenStudy (anonymous):

Okay, so step three is the error?

OpenStudy (anonymous):

Yes @Crushshroom I get x= -.167

OpenStudy (anonymous):

Step 3 shows the left side being divided by log 5. This is the change of base conversion. Notice the logs no longer have the little base 2. They are now base 10.

OpenStudy (anonymous):

I understand that. Thanks a ton ( :!!!!

OpenStudy (anonymous):

Step 3 is the error. The conversion was done incorrectly.

OpenStudy (anonymous):

This is the correct change of base formula: \(\large \log_b x = \dfrac{\log_c x}{\log_c b} \)

OpenStudy (anonymous):

That means in step 3, if the log is to be changed to base 10 from base 2, it should read: 2. \(\log_2 5 = x + 4\) is correct 3. \(\dfrac{\log 5}{\log 2} = x + 4\) is the correct step 3. Now the logs are base 10. Notice that the student did this step incorrectly bec he inverted the fraction.

OpenStudy (anonymous):

Alright, I see that. Thanks!

OpenStudy (anonymous):

You're welcome.